the resultant of two forces 3p and 2p is R. if the first force is double and the resultant is also double then find the angle between two forces

Grade:11

1 Answers

Vikas TU
7 years ago
the resultant of two forces 3p and 2p is R .i.e.
R = root(9p^2 + 4p^2 + 12p^2costhetha)
   = root(13p^2 + 12p^2*costhetha)..........................(1)
now given,
2R = root(36p^2 + 4p^2 + 24pcosthetha)
     = root(40p^2 + 24pcosthetha)...................(2)
solve these 2 eqns. and get costhetha.